Consider a two state Markov chain with one-step transition matrix on the states 1,21, , 0<p+q<2. 91-9 Show, by induction or otherwise, that the n-step transition matrix is Ptg -99 Based upon the above equation, what is lim-x P(Xn-2K-1). How about limn→x P(Xn-
Answer #1

Using induction, we first prove that the equation holds for n = 1. Then we assume that if the equation holds for n = k, then it also holds for n = k+1.

P^n = rac{1}{p+q} egin{pmatrix} q & p q & p end{pmatrix} + rac{(1 - p - q)^n}{p+q}egin{pmatrix} p &-p -q & q end{pmatrix}

As, n ightarrow infty , (1 - p - q )^n ightarrow 0 because, p + q < 2 => -1 < (1 - p - q) < 1

Thus,

linn Fn = p+q(q P) 1

So,

lim P(X,-21Xp = 1) = p/(p+ q)

lim P(X,-21Xp = 2) = p/(p+ q)

We were unable to transcribe this image

linn Fn = p+q(q P) 1

lim P(X,-21Xp = 1) = p/(p+ q)

lim P(X,-21Xp = 2) = p/(p+ q)